OpenStudy (anonymous):

marcy has a total of 100 dimes and quarters. If the total value of the coins is $14.05, how many quarters does she have?

OpenStudy (australopithecus):

x + y = 100 x(0.25) + y(0.10) = 14.05 use system of equations to solve for x and y

OpenStudy (australopithecus):

where, x = quarters and y = dimes

OpenStudy (anonymous):

put the number of quarters as \(Q\) then the number of dimes must be \(100-Q\) since they add to 100 the quarters are worth 25 cents, the dimes are worth 10 cents, and the total amount of money is \(1405\) cents set \[25Q+10(100-Q)=1405\] and solve for \(Q\)

OpenStudy (australopithecus):

team effort high five

OpenStudy (anonymous):

or @Australopithecus which shows there are a couple ways to do this there is not one "correct" way

OpenStudy (anonymous):

\[\begin{array}l\color{red}{\normalsize\text{H}}\color{orange}{\normalsize\text{I}}\color{#9c9a2e}{\normalsize\text{G}}\color{green}{\normalsize\text{H}}\color{blue}{\normalsize\text{ }}\color{purple}{\normalsize\text{F}}\color{purple}{\normalsize\text{I}}\color{red}{\normalsize\text{V}}\color{orange}{\normalsize\text{E}}\color{#9c9a2e}{\normalsize\text{}}\end{array}\]

OpenStudy (anonymous):

Okay thanks soo much! :) this really helped alot.